OpenStudy (megannicole016):

Will medal and fan!!! 1: If you are in a great deal of debt but want to save for an emergency fund, you should _______. a: save for the fund while paying down your debt b: pay down your first debt c: consider getting a second job to pay down your debt first d: consider getting a second job, and save that job's money, and then pay down your debt. 2: Your gross income is $4,520.00/month. Your deductions are FICA (7.65%),federal tax withholding (11.75%), and state tax withholding (8.5%). Your fixed expenses are 30% of your realized income. You saved 5 months' worth in an emergency fund, placing 75% in a 60-day CD at a 5.45% APR and the rest in a regular savings account at a 3.8% APR. What is the total amount of your emergency fund? How much is in the CD and savings account? How much is the total interest earned between both accounts in 60 days?

OpenStudy (megannicole016):

@Photon336 can you help?

OpenStudy (photon336):

Yeah, that's interesting. I mean just by looking at this, I can tell that you need to pay down the debt first because debt accumulates due to interest. Getting a second job would result in a second income which you could use portions of that to save for an emergency fund, while using funds from your first job to pay off your debt.

OpenStudy (megannicole016):

@Photon336 Okay, so D sounds most similar to that lol is that correct?

OpenStudy (photon336):

yeah I think d would be most reasonable

OpenStudy (megannicole016):

@Photon336 Thanks so much!! Do you know how to do the second one? That's my last one and I'm struggling..

OpenStudy (photon336):

yeah i'll take a look at it. wow that's alot of information for one question

OpenStudy (megannicole016):

Yeah no kidding lol. They save the best for last obviously -_-

OpenStudy (photon336):

Your gross income is $4,520.00/month. Your deductions are FICA (7.65%),federal tax withholding (11.75%), and state tax withholding (8.5%). Yeah so let's focus on this part first. the amount of taxes they take out. lol gotta worry about that first, because after taxes you'll have take home pay which you can use for fixed expenses. .etc

OpenStudy (megannicole016):

Okay, sounds good.

OpenStudy (megannicole016):

I seriously have no idea what to do next lol I'm so confused

OpenStudy (photon336):

well My assumption is this: we gotta subtract the FICA and taxes from the gross income.

OpenStudy (megannicole016):

Okay, let's do that then

OpenStudy (photon336):

\[(4,520)-4,520(\frac{ 7.65 }{ 100 })+4,520(\frac{ 11 }{ 100 })\]

OpenStudy (photon336):

So after taxes this is how much we have 3,457.38

OpenStudy (megannicole016):

I got 4671.42.. idk what i'm doing wrong

OpenStudy (photon336):

@megannicole016 those are taxes so you gotta subtract both from your gross income

OpenStudy (megannicole016):

792.63?

OpenStudy (photon336):

so it's 4520 sorry, so it's 842.98 in taxes.

OpenStudy (photon336):

So we have now 4520-(842.98) = $3,677

OpenStudy (megannicole016):

Okay, I think i'm following

OpenStudy (photon336):

now we take this and also consider the state tax witholdings and state tax withholding are (8.5%) so after \[3,677-3677*(\frac{ 8.5 }{ 100 }) = $3364.47\]

OpenStudy (photon336):

follow?

OpenStudy (megannicole016):

Yep, i'm following

OpenStudy (photon336):

now this: "Your fixed expenses are 30% of your realized income" what does this mean to you?

OpenStudy (megannicole016):

Does that mean you have to take 30% of $3364.47?

OpenStudy (photon336):

yeah that's what i'm thinking

OpenStudy (photon336):

30% of 3364.47 = 1009.34. that's how much is going to your fixed expenses. now 3365.47-1009.34 = $2355.13 is how much you have left over after considering taxes and fixed expenses.

OpenStudy (megannicole016):

Okay, that makes sense.

OpenStudy (photon336):

You saved 5 months' worth in an emergency fund, placing 75% in a 60-day CD at a 5.45% APR and the rest in a regular savings account at a 3.8% APR. 75% How much he put into a CD $(2355.13)*(0.75) = $1766.35 How much he put into a savings account. $588.78 so he saved 5 months worth in an emergency fund so that would end up being 5*(1766.35) = $8831.75 => amount in emergency fund. not including interest. If he saves 5 months worth in his savings account: $2943.91 => amount in savings not including interest

OpenStudy (photon336):

next and last thing we've got to do is figure out the interest. that is how much money the bank is paying you to keep your money in the savings account, and how much the CD matures in the 60 days.

OpenStudy (photon336):

usually with CDs you can't take the money out until after the period is over. do you have an idea of how you would find the interest?

OpenStudy (megannicole016):

I have no idea...

OpenStudy (photon336):

Savings account: 2,943.91 interest rate (3.8%) CD:8,831.75 (5.45%) My assumption is that the interest rate is annual so it's

OpenStudy (photon336):

usually APR is yearly. My guess is that we use the simple interest formula

OpenStudy (megannicole016):

Okay, let's try it.

OpenStudy (photon336):

\[P*R*T = (interest)\] where P is principle what you start out with R = rate in decimal form. T is the time in years. so since our period is 60 days. \[(2,943.91*(\frac{ 3.8 }{ 100 })(\frac{ 60 }{ 365 }) = $18.39| Interest~ savings\] \[(8,831.75*(\frac{ 5.45 }{ 100 })(\frac{ 60 }{ 365 }) = $79.12| Interest~ CD \] \[Total~interest = 79.12+18.39 = $97.51\]

OpenStudy (megannicole016):

Thank you thank you thank you! Is that the final answer?

OpenStudy (photon336):

yeah. my assumption was that the interest accumulates only once over the period and isn't compounded.

OpenStudy (megannicole016):

You're the best! I really appreciate it. How do I give you a medal?

OpenStudy (photon336):

just click best response and make sure you close the question afterwards